# Learning Source Biases: Multi-sourced Misspecifications and Consequences

### v2

@@ 3.3 Proof of Theorem 1
Applying the Sherman-Morrison formula to $A=\text{diag}(v_{1},\cdots,v_{M})+\mathbf{1}_{M}\mathbf{1}_{M}^{\top}$ yields
$A^{-1}=\text{diag}(\nu_{1},\cdots,\nu_{M})-\frac{1}{1+\sum_{i=1}^{M}\nu_{i}}\begin{bmatrix}\nu_{1}\\ \vdots\\ \nu_{M}\end{bmatrix}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}$
and, hence,
$\mathbf{1}_{M}^{\top}A^{-1}=\frac{1}{1+\sum_{i=1}^{M}\nu_{i}}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}.$
Substituting these results, together with $C=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}$, into Lemma 2 yields $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\overline{\Delta}$ and $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=-\overline{\Delta}$ as desired.
3.4 Discussions
When distortions of the DM’s long-run perceptions violate DOOM, they still satisfy a weaker form of DOOM (hereafter WDOOM), in that they depend only on misspecified sources and their correlations with unfamiliar sources, but nothing else. In the simple environment presented below, the comparative statics of the distortions are analogous to those described in Theorem 1.
@@ 3.4 Discussions
In the case where $X_{i,t}=b_{i}+\alpha_{i}\omega_{t}+\varepsilon_{i,t}$, $\Delta_{i}=\alpha_{i}\overline{\Delta}$ for all $i\in\mathcal{S}-\mathcal{M}$ and $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\overline{\Delta}$, where $\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\alpha_{i}\nu_{i}\Delta_{i}/(1+\sum_{i\in\mathcal{M}}\alpha_{i}^{2}\nu_{i}).$
Proposition 1 shows that when signals are correlated with the true state to varying degrees, the long-run misperceptions held by the DM will vary between sources, albeit only through $\alpha_{i}$. In addition to comparative statics with respect to $\Delta_{i}$ and $\nu_{i}$, a new insight emerges from the analysis: In the case where $\alpha_{i}>0$ $\forall i\in\mathcal{S}$ (if not, let $-X_{i}$ be the subject of analysis), increasing $\alpha_{i}$, $i\in\mathcal{S}-\mathcal{M}$ makes source $i$ more informative and therefore more susceptible to the distortionary effect of misspecified learning. In the long run, this worsens the DM’s misperception about source $i$ relative to other sources that he learns about.
The aforementioned environments satisfy a common feature, namely signals depend on each other only through the true state of the world. In the presence of additional sources of correlation, the weighting formula prescribed by Lemma 2 becomes complex very quickly, as shown by the example presented below.
Example 2.
Suppose that the period-$t$ signal generated by source $i$ is $X_{i,t}=b_{i}+\omega_{t}+\beta_{i}\theta_{t}+\varepsilon_{i,t}$, where $\theta_{t}\sim\mathcal{N}(0,1)$ represents shocks to people’s attitudes toward minorities triggered by social movements such as MeToo, and is assumed to be independent of the true state and the Gaussian errors. The presence of $\theta_{t}$ creates an additional source of correlation between the signals. The resulting covariance matrix $\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}+\beta\beta^{\top}$ is a rank-two update of a diagonal matrix, so its inverse can be obtained by iteratively applying the Sherman-Morrison formula first to $\widetilde{\Sigma}\coloneqq\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$ and then $\Sigma^{*}=\widetilde{\Sigma}+\beta\beta^{\top}$.
The algebra is brutal in general. In the case where $N=5$ and $M=3$, it can be shown that
$\begin{bmatrix}\Delta_{4}\\ \Delta_{5}\end{bmatrix}=\begin{bmatrix}\frac{P\beta_{4}+Q}{R}\\ \frac{P\beta_{5}+Q}{R}\end{bmatrix}$
where
$P=\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\Delta_{i}\right)\left(1+\sum_{i=1}^{3}\nu_{i}\right)-\left(\sum_{i=1}^{3}\nu_{i}\Delta_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right),$
$Q=\left(\sum_{i=1}^{3}\nu_{i}\Delta_{i}\right)\left(1+\sum_{i=1}^{3}\beta_{i}^{2}\nu_{i}\right)-\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\Delta_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right),$
and
$R=\left(1+\sum_{i=1}^{3}\beta_{i}^{2}\nu_{i}\right)\left(1+\sum_{i=1}^{3}\nu_{i}\right)-\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right).$
$\Delta_{4}$ and $\Delta_{5}$ depend on the correlations between familiar sources (as captured by $\beta_{i}\beta_{j}$, $i,j\in\mathcal{M}$) and those between familiar and unfamiliar sources (as captured by $\beta_{i}\beta_{j}$, $i\in\mathcal{M}$ and $j\in\mathcal{S}-\mathcal{M}$). They are independent of the correlations between unfamiliar sources (as captured by $\beta_{i}\beta_{j}$, $i,j\in\mathcal{S}-\mathcal{M}$), thus satisfying WDOOM.
3.5 Implications
We have discussed the intuitions and mathematical underpinnings of our results in previous sections. We now examine their implications for the various policy interventions used by real-world organizations to address discrimination.
Minority representation at upper echelons.
There is a large economic literature that studies the antidiscrimination effect of increasing minority representations at upper echelons. According to Bertrand and Duflo (2017), a simple mechanical mechanism may nevertheless prove to be game changing. Historically, individuals from underrepresented groups have had limited influence in decision making. To the extent that majority decision makers are biased against minority subordinates, discrimination tends to persist. The presence of minority stakeholders at the various stages of recruitment, evaluation, and promotion can serve as a counteracting force and alter the atmosphere at the top level.
Our result offers a gloomier prospect, suggesting that the damage caused by existing members of the evaluation panel may not be easily rectified by infusing new perspectives. Indeed, DOOM stipulates that the introduction of fresh, lesser-known committee members from the minority group does not alter the metric that captures the distortionary effect of misspecified learning. Diversity efforts at the top may work through alternative channels, but not through expanding the set $\mathcal{S}-\mathcal{M}$.
Reduce decision errors.
Contemporaneous rater training programs offered by numerous real-world organizations emphasize an important lesson: When making decisions that can adversely affect others (e.g., performance reviews), stakeholders should “slow down, meditate, deliberate, and follow elaborate structured procedures” so as to base decisions on facts rather than instincts (Eberhardt, 2020). Recently, Kleinberg et al. (2018) demonstrate the potential of using properly trained AI algorithms to overcome human decision errors in the context of bail decision making.
To the extent that the aforementioned interventions may increase the signal-to-noise ratios of our sources, their impacts on the outcomes of misspecified learning can be quite nuanced. For one thing, changing the signal-to-noise ratios of unfamiliar sources is ineffectual. For another, scaling the signal-to-noise ratios of all familiar sources by the same factor $\alpha>1$ increases the DM’s reliance on them when learning about the biases of unfamiliar sources, thereby amplifying the distortions in his long-run perception and behavior:
$\left|\frac{\sum_{i\in\mathcal{M}}\alpha\nu_{i}\Delta_{i}}{1+\sum_{i\in\mathcal{M}}\alpha\nu_{i}}\right|=\left|\frac{\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}}{\alpha^{-1}+\sum_{i\in\mathcal{M}}\nu_{i}}\right|>\left|\frac{\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}\right|=|\overline{\Delta}|.$
When changes occur disproportionately at different sources, the outcomes of learning may be improved due to the ambiguity of the sign of $\partial\overline{\Delta}/\partial\nu_{i}$. Yet for this approach to succeed, precise information about the misspecifications associated with each source is needed, and this knowledge is unlikely to be possessed by the DM.
Debiasing programs.
Certainly, the aforementioned interventions may operate through an alternative channel, that of correcting the misspecifications associated with familiar sources (bring $\Delta_{i}$s closer to zero). Indeed, the idea of using one’s own mental resources to debias decisions is backed by years of sociopsychological research into implicit biases, and recent evidence suggests combining it with the use of structured evaluations to prohibit decision making based on automatic, unconscious instincts (Bohnet, 2016; Greenwald and Lai, 2020).888 Outside the corporate realm, the Oakland Policy Department recently adjusted its foot pursuit policy so that officers were instructed to “step back, slow down, call for backup, and think it through” rather than to follow suspects into backyards or blind alleys. Nextdoor Neighbor, a social network for residential neighbors to communicate through, recently started asking its users to provide detailed descriptions of the suspicious activities they wish to report, because “adding frictions allows users to act based on information rather than instinct” (Eberhardt, 2020). Meanwhile, Devine et al. (2012) provide the first evidence that prejudice, stereotypes, and intergroup hostility are “bad habits” that can be broken through longitudinal interventions such as counterstereotypical thinking, priming shared identities, perspective taking, individulating, and intergroup contact. Finally, the growing literature on algorithmic fairness proposes various remedies to remove biases from algorithm design and training (Mitchell et al., 2021). In addition to interventions that can help bring $b_{i}^{*}$s closer to zero, Kalev et al. (2006) and Raji et al. (2020) stress the importance of accountability and transparency. The latter raise the DM’s awareness and consciousness and can potentially bring perceived biases ($\tilde{b}_{i}$s) closer to true biases ($b_{i}^{*}$s).
The efficacy of the aforementioned programs in addressing discrimination is a subject of active research and hot debate (more on this in Section 4.3). In case they are effective in bringing $\Delta_{i}$s closer to zero, their impacts on the DM’s belief and behavior are straightforward in the context of the baseline model. Outside the baseline model, things can become muddier, a subject to which we turn in Section 4.3.
@@ 4 Extensions
This section presents several extensions of the baseline model. We first allow the signals to depend on a common latent factor and the true state and the signals to take high-dimensional values. We then examine a problem in which both the biases of the signals and their covariance matrix are the subject of misspecified learning.
4.1 Latent factor
This section examines a variant of the baseline model in which the means of the signals can depend on a common latent factor $b_{N+1}\in\mathbb{R}$, such as employees’ productivity-enhancing investments. Redefine $b\coloneqq[b_{1},\cdots,b_{N+1}]^{\top}$ and call it fundamental. The mean vector of the signals $X\coloneqq[X_{1},\cdots,X_{N}]^{\top}$ is $\Gamma b$, where $\Gamma\coloneqq[\mathbb{I}_{N\times N}\mid\mathbf{1}_{N}]$ is the transformation matrix that operates by adding the latent factor $b_{N+1}$ to the mean of each signal.
Initially, the true value of the latent factor $b_{N+1}$, together with the biases of the sources in $\mathcal{S}-\mathcal{M}$, is unknown to the DM. Everything else is as in the baseline model. The support of DM’s prior becomes $\Theta=\{b\in\mathbb{R}^{N+1}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$.

### v3

@@ 3.3 The case of independent errors
When errors are jointly independent, their covariance matrix is simply the diagonal matrix of their variances, i.e., $\Omega^{*}=\text{diag}(v)$. Applying the Sherman-Morrison formula to the covariance matrix among familiar sources: $A=\text{diag}(v_{1},\cdots,v_{M})+\mathbf{1}_{M}\mathbf{1}_{M}^{\top}$, yields
$A^{-1}=\text{diag}(\nu_{1},\cdots,\nu_{M})-\frac{1}{1+\sum_{i=1}^{M}\nu_{i}}\begin{bmatrix}\nu_{1}\\ \vdots\\ \nu_{M}\end{bmatrix}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}.$
Substituting this finding, along with $C=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}$ as the covariance matrix between familiar and unfamiliar sources, into Theorem 1, we obtain
$\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=\frac{\sum_{i=1}^{M}\nu_{i}\Delta_{i}}{1+\sum_{i=1}^{M}\nu_{i}}\coloneqq\overline{\Delta}$
as the distortion of the DM’s prediction, and
$CA^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\overline{\Delta}$
as his misperceptions about unfamiliar sources. ∎
4 Implications
This section examines consequences of our results for various policy interventions aimed to curb discrimination, as well as their broader theoretical implications.
Minority representation at upper echelons.
There is a large economic literature that studies the anti-discrimination effect of increasing minority representations at upper echelons. According to Bertrand and Duflo (2017), a simple mechanical mechanism may nevertheless prove to be game changing: Historically, individuals from underrepresented groups have had limited influence in decision making. To the extent that majority decision makers are biased against minority subordinates, discrimination tends to persist. The presence of minority stakeholders at the various stages of recruitment, evaluation, and promotion can serve as a counteracting force and alter the atmosphere at the top level.
Our result offers a gloomier prospect, suggesting that the damage caused by existing members of the evaluation panel may not be easily rectified by infusing new perspectives. Indeed, DOOM stipulates that minority employees are underpaid to the same degree, with or without the introduction of fresh, lesser-known committee members from the minority group to the evaluation panel. Diversity efforts at the top may work through alternative channels, but not through expanding the set of lesser-known sources.
Reduce decision errors.
Contemporaneous rater training programs offered by numerous real-world organizations to curb discrimination emphasize an important lesson: When making decisions that can adversely affect others (e.g., performance reviews), stakeholders should “slow down, meditate, deliberate, and follow elaborate structured procedures” so as to base decisions on facts rather than instincts (Eberhardt, 2020). Recently, Kleinberg et al. (2018) demonstrate the potential of using properly trained AI algorithms to overcome human decision errors in the context of bail decision making.
To the extent that the aforementioned interventions may increase the precision of our sources, their impacts on the outcomes of misspecified learning can be quite nuanced. For one thing, changing the signal-to-noise ratios of unfamiliar sources is ineffectual because of DOOM. For another, the success of changing the signal-to-noise ratios of familiar sources hinges on the nature of misspecifications. In the case of independent errors, perturbing the signal-to-noise ratios of familiar sources by $(d\nu_{i})_{i\in\mathcal{M}}$ changes $\overline{\Delta}$ by
$\frac{\sum_{i\in\mathcal{M}}d\nu_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}\left(\sum_{i\in\mathcal{M}}\Delta_{i}\kappa_{i}-\overline{\Delta}\right),\text{ where }\kappa_{i}\coloneqq\frac{d\nu_{i}}{\sum_{j\in\mathcal{M}}d\nu_{j}}.$
The impact depends not only on the total perturbation $\sum_{i\in\mathcal{M}}d\nu_{i}$, but also on the relative perturbations $\kappa_{i}$ of the various sources, weighted by their respective misspecifications $\Delta_{i}$. For this approach to succeed, one needs precise information about the misspecifications to make astute choices about $\kappa_{i}$. [^8] In particular, naively scaling the signal-to-noise ratios of all misspecified sources by the same factor $\alpha>1$ results in $d\nu_{i}=(\alpha-1)\nu_{i}$ and $\kappa_{i}=\nu_{i}/\sum_{j\in\mathcal{M}}\nu_{j}$. The change in $\overline{\Delta}$ equals $\frac{(\alpha-1)\overline{\Delta}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}$, hence the distortionary effect of misspecified learning is amplified by the intervention. In practice, this knowledge is unlikely to be possessed by the DM.
Debiasing programs.
Certainly, the aforementioned interventions may operate through an alternative channel, that of (partially) correcting the misspecifications associated with familiar sources. Indeed, the idea of using one’s own mental resources to debias decisions is backed by years of sociopsychological research into automatic, unconscious biases (Greenwald and Lai, 2020).999 Outside the corporate realm, the Oakland Police Department recently adjusted its foot pursuit policy so that officers were instructed to “step back, slow down, call for backup, and think it through,” rather than follow suspects into backyards or blind alleys. Nextdoor, a social network for residential neighbors to communicate through, recently started asking its users to provide detailed descriptions of the suspicious activities they wish to report, because “adding frictions allows users to act based on information rather than instinct” (Eberhardt, 2020). Meanwhile, Devine et al. (2012) provide the first evidence that prejudice, stereotypes, and intergroup hostility are “bad habits” that can be broken through longitudinal interventions such as counterstereotypical thinking, priming shared identities, and intergroup contact. Finally, the growing literature on algorithmic fairness proposes various remedies to remove biases from algorithm design and training (Mitchell et al., 2021).
While the above interventions aim to bring true source biases $b_{i}^{*}$ closer to zero, another approach is to align perceived biases $\tilde{b}_{i}$ with true biases. This can be achieved by enhancing accountability and transparency within the organization (Kalev et al., 2006; Raji et al., 2020), therefore revealing the biases of lower-ranked raters to the DM. Alternatively, the DM can consult third parties who, without any preconception, estimate the biases of familiar and unfamiliar sources simultaneously. [^10] This is the essence of the famous Durbin-Wu-Hausman test that econometricians use to detect model misspecification (Patrick, 2020). These efforts can raise the DM’s awareness of potential misspecification, (hopefully) prompting a revision of perceptions about familiar sources.
The efficacy of the aforementioned programs in addressing discrimination is a subject of active research and hot debate (more on this in Section 5.1). In case they are effective in bringing misspecifications closer to zero, their impacts on the DM’s beliefs and prediction are straightforward in the context of the baseline model. Outside the baseline model, things can become muddier, as we shall demonstrate in Section 5.1.
Alternative decision situations.
Consider a common scenario in finance, consulting, and service industries, where a DM informs his client of the true state $\omega$ using a linear model $X=b+\omega\mathbf{1}_{N}+\epsilon$. Suppose $\omega$ and $\epsilon$ have mean zero but may not follow normal distributions, and $\omega$ has unit variance. When estimating his model, the DM fixes the mean of $X_{i}$, $i\in\mathcal{M}$, to $\tilde{b}_{i}$, consistent with industry standards or assumptions he deems appropriate. In case these assumptions deviate from reality, the DM’s estimator — which minimizes the estimation squared-error — becomes the constrained least squares estimator defined in Footnote 7. Based on this, the DM makes a linear prediction to minimize the prediction squared-error. Formally, a linear predictor for $\omega$ in $X$ has the form $\beta^{\top}(X-b)$. The best linear predictor, BLP, is the one that minimizes prediction squared-error $\mathbb{E}[\left(\omega-\beta^{\top}(X-b)\right)^{2}]$.
@@ 4 Implications
In situations described above, the distortion in BLP is as in Theorem 1 and in particular satisfies DOOM.
Identify misspecified sources.
DOOM can be used to identify misspecified sources. In the context of the leading example, suppose that we, as researchers, can alter the covariance matrix of the signals through informational interventions, such as rater training, as discussed in Section 3.3.111111Similar types of informational interventions have been proposed by Bohren et al. (2023) as means to distinguish discrimination based on erroneous beliefs from traditional taste-based or rational statistical discrimination. However, the issue of misspecified learning of the biases of novel sources and its relationship between the persistent wage gap between majorities and minorities has eluded previous researchers. DOOM stipulates that the wage gap between majority employees — who are exempted from unconscious biases — and minority employees — who are subject to unconscious biases — should vary only with the covariances between misspecified sources, but not the covariances between novel sources or those between novel and misspecified sources.
Alternative forms of misspecification.
We model misspecification as degenerate, wrong, prior beliefs about familiar sources, consistent with the misspecified Bayesian learning literature. However, the qualitative nature of our results remains valid if the support of the prior is an unbounded convex polyhedron in $\mathbb{R}^{N}$. [^12] The assumption of unboundedness ensures that the limiting posterior beliefs concentrate on the set of KL minimizers. An example would be $\Theta=\times_{i=1}^{N}[\underline{b}_{i},\overline{b}_{i}]$, where $[\underline{b}_{i},\overline{b}_{i}]$ is unbounded for some $i$. Under this more general assumption, the KL minimization problem remains a convex optimization problem, solvable using the Lagrangian method. Let $\mathcal{M}$ denote the dimensions involved in binding linear constraints and call $\mathcal{M}$ the endogenous set of familiar sources. Given $\mathcal{M}$, all previous results, including DOOM, remain valid.
@@ 5 Extensions
This section examines two extensions of the baseline model, one in which the biases of the signals and their covariance matrix are both the subject of misspecified learning, and another in which the true state and the signals can take high-dimensional values.
5.1 Unknown covariance matrix
So far we have assumed that the covariance matrix of the signals is known to the DM from the outset. In environments such as large organizations with numerous divisions and hierarchies, a more natural assumption seems to be that the covariance matrix is also a subject of misspecified learning. We consider this possibility in this section, with a focus on whether the additional source of misspecified learning increases the bias in the DM’s beliefs and prediction or ameliorates his bias.
Formally, fix the covariance matrix of the errors to any $\Omega^{*}\in\mathbb{R}^{N\times N}$ that is symmetric and positive definite (SPD). Fix any $\underline{\lambda}\in(0,\min\{1,\lambda_{\min}(\Omega^{*})\})$, and define $\Lambda$ as the set of $N\times N$ SPD matrices whose minimum eigenvalues are at least $\underline{\lambda}$. Initially, the DM holds prior beliefs about the distribution of the signals. The prior is correct about signals equalling their biases plus the true state and state-independent errors. However, it misspecifies the biases of some sources as in Section 2, while allowing the covariance matrix of the errors to be anything in $\Lambda$. The support of the prior, again denoted by $\Theta$, is
Importantly, $\Theta$ contains the true covariance matrix of the signals $\Sigma^{*}=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. This means that while the DM is initially uncertain about the covariance matrix, he never holds any misspecification against it.
We characterize outcomes of misspecified learning in two steps, beginning with the following lemma. [^13] In the Online Appendix, we provide an alternative, frequentist, foundation for these limiting beliefs.
Lemma 2.
In the environment considered in this section, the DM’s long-run beliefs concentrate on the set of distributions that minimize the KL divergence (1), subject to the constraint that $(b,\Sigma)\in\Theta$.
Let $\mathcal{N}(\hat{b},\widehat{\Sigma})$ denote the KL minimizer, which is shown to be unique. Define the DM’s misperception of source biases, as well as the distortion of his long-run prediction about the true state, as before. The next theorem pinpoints the difference between these quantities and their analogs in the baseline model.
Theorem 2.
In the environment considered in this section,
  1. (i)
    The DM’s misperception about source biases $\Delta$ is the same as in the baseline model, while his limiting belief about the covariance matrix of the signals is now given by $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$.
  2. (ii)
    The expected distortion of the DM’s prediction, evaluated under the true signal distribution, equals its counterpart in the baseline model, discounted by $\delta\coloneqq(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}\in(0,1)$.
  3. (iii)
    The result in Part (ii) satisfies DOOM. When true errors are jointly independent, it equals $-\delta\overline{\Delta}$, and its comparative statics with respect to $\nu_{i}$ and $\Delta_{i}$, $i\in\mathcal{M}$ are generally ambiguous.
Part (i) of Theorem 2 shows that on the one hand, misspecified learning of the covariance matrix does not interfere with that of source biases, leaving the DM’s long-run misperceptions about the latter unaffected (more on this later). On the other hand, the inconsistency caused by misspecifications makes signals appear noisier to the DM than they truly are (in the Loewner order), in that the perceived covariance matrix equals its true counterpart, plus a rank-one update $\Delta\Delta^{\top}$ that is positive definite. Consequently, when making predictions about the true state, the DM relies less on signal realizations compared to when he knows the true covariance matrix from the outset. As shown in Part (ii) of Theorem 2, this mitigates the impact of misspecified learning on his behavior, resulting in a less biased prediction on average.
Turning to the comparative statics in Part (iii), what is new here is the ambiguity associated with correcting misspecifications, whose impact is unambiguously positive in the baseline model (recall Part (ii) of Proposition 1). The ambiguity arises from compound misspecified learning: Correcting the misspecifications associated with familiar sources makes the environment appear less noisy to the DM than before. This reduces the bias mitigation effect discussed above and, as confirmed by numerical analysis, could undo the benefit of misspecification corrections.
Our finding sheds light on the behavioral and welfare consequences of the debiasing programs introduced in Section 4. Years of economic and psychological research have yielded mixed insights into the efficacy of these programs. For example, while the success of using one’s own mental resources to intercept implicit biases has been documented in prominent case studies (recall Footnote 9), a recent meta-analysis of organizations that have implemented similar programs reveals a mixed pattern. This has led Greenwald and Lai (2020) to conclude that “Convincing evidence for the effectiveness of these strategies is not yet available in peer-reviewed publications.” Even worse, efforts to suppress stereotypes or to reduce intergroup animosity may backfire (Kulik et al., 2000; Kalev et al., 2006). We take no stand on this debate and instead caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$ closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect.
Proof sketch.
We outline the proof of Theorem 2, focusing on why compound misspecified learning does not further distort the DM’s perception about source biases, as well as the bias mitigation effect.
A closer inspection of the KL minimization problem reveals that with compound misspecified learning, the DM’s perception about source biases is simply the solution to problem (2), given his limiting belief about the covariance matrix $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$. What remains to be verified is that the distortion in the covariance matrix, $\Delta\Delta^{\top}$, does not affect the solution to problem (2). Define $\widehat{A}$ as the submatrix of $\widehat{\Sigma}$ restricted to familiar sources, and $\widehat{C}$ as the submatrix capturing the covariances between familiar and unfamiliar sources. Our claim boils down to showing that $\widehat{C}\widehat{A}^{-1}\widetilde{\Delta}=CA^{-1}\widetilde{\Delta}$.
By definition, $\widehat{A}=A+\widetilde{\Delta}\widetilde{\Delta}^{\top}$ and $\widehat{C}=C+\widehat{\Delta}\widetilde{\Delta}^{\top}$, where $A$ and $C$ are the true counterparts of $\widehat{A}$ and $\widehat{C}$, $\widetilde{\Delta}$ is the vector of the misspecifications associated with familiar sources, and $\widehat{\Delta}$ is the vector of optimized misperceptions about unfamiliar sources in the baseline model. Expanding $\widehat{C}\widehat{A}\widetilde{\Delta}-CA\widetilde{\Delta}$ yields
$\widehat{C}\widehat{A}^{-1}\widetilde{\Delta}-CA^{-1}\widetilde{\Delta}$ $=(C+\widehat{\Delta}\widetilde{\Delta}^{\top})(A^{-1}-\frac{A^{-1}\widetilde{\Delta}\widetilde{\Delta}^{\top}A^{-1}}{1+\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}})\widetilde{\Delta}-CA^{-1}\widetilde{\Delta}$
$=\underbrace{-\frac{CA^{-1}\widetilde{\Delta}\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}{1+\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}}_{\text{(I)}}+\underbrace{\frac{\widehat{\Delta}\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}{1+\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}}_{\text{(II)}}=0,$
where the first equality uses the Sherman-Morrison formula to calculate $\widehat{A}^{-1}$, and the last equality uses $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ (recall Theorem 1). Intuitively, distortions in the covariance matrix affect the DM’s perception about source biases through two channels. The first arises from familiar sources and is captured by $\widehat{A}^{-1}-A^{-1}=-\frac{A^{-1}\widetilde{\Delta}\widetilde{\Delta}^{\top}A^{-1}}{1+\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}$. The second concerns the covariances between familiar and unfamiliar sources and is captured by $\widehat{C}-C=\widehat{\Delta}\widetilde{\Delta}^{\top}$. To assess the overall impact, we project the first effect onto unfamiliar dimensions using $C$. Since this same projection is used to derive the optimal misperceptions about unfamiliar sources in the baseline model, i.e., $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$, the two effects must cancel each other.
Turning to the bias mitigation effect, we decompose the distortion in the DM’s long-run prediction as follows:
Effect (I) in the above expression is due to the misspecified learning of source biases and is the focus of the baseline model. Effect (II) arises from the misspecified learning of the covariance matrix and has zero mean under the true signal distribution. Effect (III) captures the compound effect of misspecified learning. It equals